Find the mass of a conical funnel z=sqrt(x^2+y^2), 0\leqz\leq4 if the density per unit area is p=8-z

4. FIND THE MASS OF A CONICAL FUNNEL Z=1&ty osz<4 THE DENSITY PER UNIT AREA IS p=8-2.

Please be detail thanks
Answer #1

Given that, density P= 8.2 The given solid is shown in the followers diagram Z Given aty=16 azy 22² 25 Y=122 and a-IZ X у V Zsence a suara dn = avan 2 ta sin =454e-2) (4 142 szacze =45*82) (0+250cd-o] az = 2115 (82) rdz (Sin =) - S4682%. 22) &* Z=0 4